OpenStudy (mikaylas):

(n + 1) = –10f(n). If f(1) = 1, what is f(3

OpenStudy (bvfiywerhfhyei):

wed udiwedn weuhduiw ediwheiudhiuweud weudhuwehduhweudhewud wuhduiweh duiew dewh duew udhewuihdweudhweiud wedhiweiwedewdhwedwe weudhuewduwehduwhu w we we idweh udhewudhuiwe iuwe

OpenStudy (saksh1998):

just substitute the value of n as 3 fan and medal

OpenStudy (bvfiywerhfhyei):

\[\le \sqrt{_{\sqrt[\prime _{^{\cot \tan \sec \exp \tan^{-1} \left( \left\{ \left(\begin{matrix}\left[\begin{matrix}math & ? \\ ? & ?\end{matrix}\right] \\ ?\end{matrix}\right) \right\} \right)}}^{?}]{?}}}\]

OpenStudy (bvfiywerhfhyei):

dewedwedewdwedw

OpenStudy (bvfiywerhfhyei):

ewdwedwedewdwed

OpenStudy (bvfiywerhfhyei):

chicken

OpenStudy (bvfiywerhfhyei):

wedwed

OpenStudy (bvfiywerhfhyei):

wed

OpenStudy (bvfiywerhfhyei):

wedwedwedewdewdwed